how do you graph the equation x-4y=-4
Answer by ewatrrr(24785)


Hi
Note: the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
x-4y = -4 |solving for y to put into the slope-intercept form
y = (1/4)x + 1 |slope m = (1/4), y-intercept is Pt(0,1)
Graph by plotting Pt(0,1) and using slope to determine other points on line
Note: as y increases/decreases by 1, x increases/decreases by 4
For ex: (4,2) and (-4,0) are on the line